I need help solving this question: A cylinder tank has a diameter fo 20 inches and a height of 30 inches. In it is a solid cylinder that has a diameter fo 12 inches and a height fo 30 inches. If the tank is filled with water, what percent of the original tank can be used for water?

Answer by nerdybill(2712)

I need help solving this question: A cylinder tank has a diameter fo 20 inches and a height of 30 inches. In it is a solid cylinder that has a diameter fo 12 inches and a height fo 30 inches. If the tank is filled with water, what percent of the original tank can be used for water?
.
V = (pi)r^2h
.
Volume of original tank =
(3.14)(10^2)(30)
(3.14)(100)(30)
(3.14)(3000)
9420 cubic inches
.
Volume of solid cylinder =
(3.14)(6^2)(30)
(3.14)(36)(30)
(3.14)(1080)
3391.2 cubic inches
.
percent of the original tank can be used for water:
3391.2/9420 * 100
= 0.36 * 100
= 36%